Req 1 Req 2 and 3 Calculate the companys total estimated direct labor cost for each quarter of the the upcoming fiscal yearReq 1 Req 2 and 3 Calculate the companys total estimated manufacturing overhead cost and the cash disbursements for manufactThe Production Department of Hruska Corporation has submitted the following forecast of units to be produced by quarter for the upcoming fiscal year: 1st Quarter 2nd Quarter 3rd Quarter 4th Quarter Units to be produced 10,200 9,200 11,200 12,200 Each unit requires 0.25 direct labor-hours and direct laborers are paid $11.00 per hour. In addition, the variable manufacturing overhead rate is $1.60 per direct labor-hour. The fixed manufacturing overhead is $82,000 per quarter. The only noncash element of manufacturing overhead is depreciation, which is $22,000 per quarter. Required: 1. Calculate the company’s total estimated direct labor cost for each quarter of the upcoming fiscal year and for the year as a whole. Assume that the direct labor workforce is adjusted each quarter to match the number of hours required to produce the estimated number of units produced. 2&3. Calculate the company’s total estimated manufacturing overhead cost and the cash disbursements for manufacturing overhead for each quarter of the upcoming fiscal year and for the year as a whole.
Answer #1

Answer-1)-

HRUSKA CORPORATION
TOTAL DIRECT LABOR COSTS
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Direct labor costs 28050 25300 30800 33550 117700

Explanation-

HRUSKA CORPORATION
TOTAL DIRECT LABOR COSTS
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Units to be produced (a) 10200 9200 11200 12200 42800
Each units requires hours (b) 0.25 0.25 0.25 0.25 0.25
Budgeted direct labor hours (c=a*b) 2550 2300 2800 3050 10700
Direct labor rate per hour (d) $11.00 $11.00 $11.00 $11.00 $11.00
Direct labor costs (e=c*d) $ 28050 25300 30800 33550 117700

2)-

HRUSKA CORPORATION
TOTAL DIRECT LABOR COSTS
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Total manufacturing overhead $ 86080 85680 86480 86880 345120

Explanation-

HRUSKA CORPORATION
Manufacturing Overhead budget
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Units to be produced (a) 10200 9200 11200 12200 42800
Each units requires hours (b) 0.25 0.25 0.25 0.25 0.25
Budgeted direct labor hours (c=a*b) 2550 2300 2800 3050 10700
Variable manufacturing overhead rate per DLH (d) $1.60 $1.60 $1.60 $1.60 $1.60
Variable manufacturing overhead (e=c*d) $ 4080 3680 4480 4880 17120
Add:- Fixed manufacturing overhead $ 82000 82000 82000 82000 328000
Total manufacturing overhead $ 86080 85680 86480 86880 345120

3)-

HRUSKA CORPORATION
Manufacturing Overhead budget
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Cash disbursements for manufacturing overhead $ 64080 63680 64480 64880 257120

Explanation-

HRUSKA CORPORATION
Manufacturing Overhead budget
Particulars First Quarter Second Quarter Third Quarter First Quarter Year
Units to be produced (a) 10200 9200 11200 12200 42800
Each units requires hours (b) 0.25 0.25 0.25 0.25 0.25
Budgeted direct labor hours (c=a*b) 2550 2300 2800 3050 10700
Variable manufacturing overhead rate per DLH (d) $1.60 $1.60 $1.60 $1.60 $1.60
Variable manufacturing overhead (e=c*d) $ 4080 3680 4480 4880 17120
Add:- Fixed manufacturing overhead $ 82000 82000 82000 82000 328000
Total manufacturing overhead $ 86080 85680 86480 86880 345120
Less:- Depreciation $ 22000 22000 22000 22000 88000
Cash disbursements for manufacturing overhead $ 64080 63680 64480 64880 257120